Using the information below, answer the following questions

a. Assume a private closed economy, find the multiplier & equilibrium GDP

b. Graph the data provided in table above, and show the equilibrium point graphically. Calculate changes in

unplanned inventory at each level of GDP.

c. Discuss mathematically different conditions that prevail at the equilibrium point for a private closed economy

d. Assume an open economy, find the multiplier & equilibrium GDP

e. Use the same graph in (b) and show the change in equilibrium as a result of switch from closed to open economy

f. Now assume that instead of Imports = 20 , the new import function M=10 +0.1Y , how would the equilibrium

point change


1
Expert's answer
2021-03-18T15:22:28-0400

a. In a private closed economy the multiplier & equilibrium GDP are:

Y = C + I + G,

"m = \frac{1} {1 - c(1 - t)}"

b. The equilibrium point can be shown as intersection of AD and AS curves.

c. At the equilibrium point for a private closed economy Y = C + I + G, C = C0 + c(Y - T), T = tY.

d. In an open economy the multiplier and equilibrium GDP are:

Y = C + I + G + NX,

"m = \frac{1} {1 - c(1 - t) + m}"

e. The change in equilibrium as a result of switch from closed to open economy depends on the amount of net exports (NX).

f. If instead of Imports = 20 , the new import function M = 10 + 0.1Y, then the equilibrium would change depending on the amount of Y.